OpenStudy (anonymous):

Evaluate 2x − y when x = 3 and y = −2

OpenStudy (ajprincess):

jst substitute the values of x and y in 2x − y

OpenStudy (anonymous):

is it 4? or negative 4

OpenStudy (anonymous):

I get confused with the signs

OpenStudy (ajprincess):

2x-y=2(3)-(-2)=6+2

OpenStudy (ajprincess):

(-)*(-)=(+) (+)(-)=(-) (-)(+)=(-) (+)(+)=(+)

OpenStudy (ajprincess):

is that clear nw?

OpenStudy (anonymous):

Yes.